Hi,

Please answer clearly with details.

Thank you so much for your help!!!

in a cylinder at a constant pressure of 2.30 105 is cooled and compressed from 1.70 m3 to 1.20 m3. The internal energy of the gas decreases by 1.40 x 105 J. (a) Find the work done by the gas. (b) Find the amount of the heat that flowed into or out of the gas. C) State the direction (inward or outward) of the flow. Explain. (answer: a) -1.15 x 105 J b) 2.25 105 c) heat of the
Answer #1

Pressure P = 2.3* 10 ^ 5 Pa

Initial volume V = 1.7 m^ 3

Final volume V ‘ = 1.2 m ^ 3

Decrease in internal energy dU = -1.4 * 10^5 J

Work done by the gas W = P ( V ‘ – V )

= -115000 J

(b). heat Q = dU + W

= -255000 J

Q is negative .So, heat flow out of the gas

(c). Gas is not ideal.

Since there is an change in internal energy